OpenStudy (anonymous):

Realllyyy need help with this question... The Cool Company determines that the supply function for its basic air conditioning unit is S(p) = 30 + 0.006p3 and that its demand function is D(p) = 150 - 0.12p2, where p is the price. Determine the price for which the supply equals the demand.

OpenStudy (anonymous):

set the supply function = demand function and solve for p.

OpenStudy (whpalmer4):

To find the price where supply equals demand, set S(p) = D(p) and solve for p. \[30 + 0.006p^3 = 150 - 0.12p^2\]